# Deep Class-Incremental Learning: A Survey

### v1

@@ 3.2.1 Dynamic Networks
Recently, Vision Transformer (ViT) [24] has attracted the attention of the computer vision community, and many works tend to design CIL learners using ViT as the backbone. DyTox [23] is the first work to explore ViT in CIL, which finds that model expansion in ViT is much easier than in convolutional networks. In DyTox, only task tokens are expanded for each new task, which requires much less memory than saving the whole backbone. Similarly, L2P [72] and DualPrompt [73] explore how to build CIL learners with pre-trained ViT. They borrow ideas from Visual Prompt Learning (VPT) [145] to incrementally finetune the model with prompts. In L2P, the pre-trained ViT is frozen during the learning process, and the model only optimizes the prompts to fit new patterns. The prompt pool is defined as:
$\mathbf{P}=\left\{P_{1},P_{2},\cdots,P_{M}\right\}\,$, (11)
where $M$ is the total number of prompts, $P_{i}\in\mathbb{R}^{L_{p}\times d}$ is a single prompt with token length $L_{p}$ and the same embedding size $d$ as the instance embedding $\phi({\bf x})$. The prompts are organized as key-value pairs — each instance selects the most similar prompts in the prompt pool via KNN search. It obtains instance-specific predictions by adapting the input embeddings as:
${{\bf x}}_{p}=\left[P_{s_{1}};\cdots;P_{s_{N}};\phi({\bf x})\right],\quad 1\leq N\leq M\,.$ (12)
$P_{s_{j}}$ are selected prompts via KNN search. The adapted embeddings are then fed into the self-attention layers [146] to obtain instance-specific representations. CODA-Prompt [74] extends the prompt search with the attention mechanism. Apart from pre-trained ViT, S-Prompt [75] utilizes the pre-trained language-vision model CLIP [147] for incremental learning, which simultaneously learns language prompts and visual prompts to boost representative embeddings. [76] extends CLIP with spatial and temporal prompts for incremental video classification.
3.2.2 Parameter Regularization
@@
Dynamic networks seek to adjust model capacity with data evolves. However, if the model structure is fixed and unchangeable, how to adjust the plasticity to resist catastrophic forgetting? Parameter regularization methods consider that the contribution of each parameter to the task is not equal. Hence, they seek to evaluate the importance of each parameter to the network and keep the important ones static to maintain former knowledge.
@@
The parameter-wise regularization term is calculated based on two parts. $\theta_{k}^{b-1}$ denotes the $k$-th parameter after learning last task $\mathcal{D}^{b-1}$. Hence, $(\theta_{k}^{b-1}-\theta_{k})^{2}$ represents the parameter drift from the last stage, and $\Omega_{k}$ weighs it to make sure important parameters do not shift away from the last stage. Since the model at the last stage represents the ‘old’ knowledge, consolidating important parameters can fix the knowledge from being forgotten.
@@
Eq. 13 depicts a way to penalize essential parameters, and there are different ways to calculate the importance matrix $\Omega$. In EWC, Fisher information matrix [148] is adopted to estimate $\Omega$. However, the importance calculation in EWC is conducted at the end of each task, which ignores the optimization dynamics along the model training trajectory. To this end, SI [78] proposes to estimate $\Omega$ in an online manner and weigh the importance via its contribution to loss decay. RWalk [39] combines these importance estimation techniques. [79, 80] resort to an extra unlabeled dataset for online evaluation. IMM [81] finds a maximum of the mixture of Gaussian posteriors with the estimated Fisher information matrix. IADM [82] and CE-IDM [83] analyze the capacity and sustainability of different layers in class-incremental learning and find that different layers have different characteristics in CIL. In detail, shallow layers converge faster but have limited representation ability. By contrast, deep layers converge slowly while having powerful discrimination abilities. Hence, IADM augments EWC with an ensemble of different layers and learns layer-wise importance matrix in an online manner. K-FAC [84] extends the Fisher information matrix approximation with the Kronecker factorization technique.
3.2.3 Discussions about Model-Centric Methods
Model-centric methods either design model structure adjustment techniques or build generalization terms to resist forgetting. Apart from these groups, there are works addressing network masks to divide a large network into sub-networks for each task [149, 150, 151, 152]. However, deciding the activation of a specific sub-network requires the task identifier or learning extra task classifiers. On the other hand, several works [153, 154, 155] propose to design specific modules for incremental new tasks, e.g., adapters [156]. However, manually handcrafting these modules requires heuristic designs or task-specific priors.
Learning dynamic networks, especially backbone expansion methods, have achieved state-of-the-art performance in recent years [22]. However, it often requires expandable memory budgets, which is unsuitable for incremental learning on edge devices. Secondly, pre-trained models are needed for the prompt expansion methods. However, a pre-trained model is not always available for some specific downstream tasks, e.g., face recognition [157], speech recognition [158], etc. Therefore, how to get rid of the dependence on pre-trained models is essential for these methods in real-world applications.
@@ 3.2.3 Discussions about Model-Centric Methods
On the other hand, parameter regularization methods have attracted the attention of the community in the early years. However, estimating parameter importance requires saving the matrix with the same scale as the model. It faces the same risk as dynamic networks that the memory budget is linearly increasing. On the other hand, the importance matrix may conflict at different incremental stages [92], making it hard to optimize the model.
Figure 5: Illustration of knowledge distillation methods in class-incremental learning. Left: Logit distillation (e.g., iCaRL [32]) aligns the model outputs to make the old model and new model share the same semantic relationship. Middle: Feature distillation (e.g., UCIR [93]) aligns the features produced by the old and new models to make sure the new model does not forget old features. Right: Relational distillation (e.g., R-DFCIL [101]) resorts to structural inputs, e.g., triples, and aligns the input relationship of the old and new model.
3.3 Algorithm-Centric Class-Incremental Learning
Algorithm-centric CIL methods focus on designing algorithms to maintain the model’s knowledge in former tasks. For example, the core idea of knowledge distillation-based CIL methods is to build the mapping between old and new models and reflect the characteristics of the old model in the updating process. On the other hand, model rectify-based methods aim to discover and reduce the bias in incremental models.
3.3.1 Knowledge Distillation
Training data is evolving in the incremental learning process, requiring tuning the model sequentially. We can denote the model after the previous stage $f^{b-1}$ as the ‘old model’ and the current updating model $f$ as the ‘new model.’ Assuming the old model is a good classifier for all the seen classes in $\mathcal{Y}_{b-1}$, how can we utilize it to resist forgetting in the new model?
@@ 3.3.1 Knowledge Distillation
To enable the old model to assist the new model, an intuitive way utilizes the concept proposed in [159, 160, 144], i.e., knowledge distillation (KD). KD enables the knowledge transfer from a teacher model to the student model, with which we can teach the new model not to forget. There are several ways to build the distillation relationship, and we divide these KD-based methods into three subgroups, i.e., logit distillation, feature distillation, and relational distillation.
@@
LwF [85, 86] is the first success to apply knowledge distillation into CIL. Similar to Eq. 13, it builds the regularization term via knowledge distillation to resist forgetting:
@@
where the old model ${f}^{b-1}$ is frozen during updating. The regularization term builds the mapping between the old and new models by forcing the predicted probability among old classes to be the same. Given a specific input $\bf x$}, the output probability of the $k$-th class reveals the semantic similarity of the input to this class. Hence, Eq. 14 forces the semantic relationship of the old model and new model to be the same and resists forgetting. iCaRL [32] extends LwF with the exemplar set, which helps to further recall former knowledge during incremental learning. Additionally, it drops the fully-connected layers and follows [161] to utilize nearest-mean-of-exemplars during inference. Eq. 14 strikes a trade-off between old and new classes, where the former part aims to learn new classes and the latter one maintains old knowledge. Since the number of old and new classes may differ in different incremental stages, BiC [87] extends Eq 14 by introducing a dynamic trade-off term:
@@
Eq. 15 forces the features extracted by the new embedding module to be the same as the old one, which is a stronger regularization than Eq. 14. Several works follow it to utilize feature distillation in CIL [94, 95], while others address distilling other products. LwM [96] suggests penalizing the changes in classifiers’ attention maps to resist forgetting. AFC [97] conducts the distillation considering the importance of different feature maps. PODNet [98] minimizes the difference of the pooled intermediate features in the height and width directions instead of performing element-wise distillations. CVIC [162] decouples the distillation term into spatial and temporal features for video classification. DDE [99] distills the causal effect from the old training to preserve the old knowledge. GeoDL [100] proposes to conduct distillation based on the projection of two sets of features from old and new models.
@@
3.3.2 Model Rectify
@@
The first method addressing the bias of the CIL model is UCIR [93]. It finds that the weight norm of new classes is significantly larger than old ones, and the model tends to predict instances as the new classes with larger weights. Hence, UCIR proposes to utilize a cosine classifier to avoid the influence of biased classifiers: $f({\bf x})=\frac{W}{\lVert W\rVert}^{\top}\frac{\phi({\bf x})}{\lVert\phi({\bf x})\rVert}$. Hence, the weight norm will not influence model predictions in incremental learning. WA [115] further normalizes the weight after every optimization step. It also introduces weight clipping to ensure the predicted probability is proportionate to classifier weights. SS-IL [116] explains the reason for weight drifting, which is caused by the imbalance phenomena between old and new instances. Since the number of new class instances is much more than that of old ones, optimizing the model with cross-entropy loss will increase the weight of new classes and decrease old ones. Hence, SS-IL suggests separated softmax operation and task-wise knowledge distillation to alleviate the influence of imbalanced data. RPC [117] claims that the classifiers of all classes can be pre-allocated and fixed. This makes it impossible for classifiers to be biased toward new classes.
@@
Note that only the logits for new classes ($k\in{Y}_{b}$) are rectified after each incremental task. BiC separates an extra validation set from the exemplar set, i.e., $\mathcal{E}=\mathcal{E}_{train}\cup\mathcal{E}_{val}$, and uses the validation set to tune the rectification layer. On the other hand, IL2M [113] suggests re-scaling the outputs with historical statistics. Suppose an instance is predicted as a new class. In that case, the logits will be re-scaled to ensure the predictions of old and new classes follow the same distribution. AANets [114] learns to adaptively aggregate the stable and plastic blocks and adjust the predictions.
@@
Lastly, since the embedding module is sequentially updated in CIL, some works try to rectify the biased representations of incremental models. For example, SDC [107] utilizes a nearest-mean-of-exemplars classifier, which calculates the class centers and assigns instances to the nearest class center. However, since the embedding is incrementally updated, the class centers calculated in the former stage may suffer a drift in the next stage, making the classification results unreliable. Since old class instances are not available in the current stage, SDC aims to calibrate the class centers of old classes with the drift of new classes. CwD [108] analyzes the differences of embeddings among the CIL model and oracle model and finds that the embeddings of the oracle model scatter more uniformly. It aims to make the CIL model similar to the oracle by enforcing the eigenvalues to be close. ConFiT [109] relieves the feature drift from the middle layers. Other works address the model weight rectification. CCLL [110] aims to calibrate the activation maps of old models during incremental learning. RKR [111] proposes to rectify the convolutional weights of old models when learning new tasks. Recently, FACT [106] depicts a new training paradigm for CIL, namely forward compatible training. Since the embedding space is endlessly adjusted for new classes, FACT proposes to pre-assign the embedding space for new classes to relieve the burden of embedding tuning.
3.3.3 Discussions about Algorithm-Centric Methods
Algorithm-centric methods design task-specific algorithms to resist forgetting. Knowledge distillation-based methods have been widely applied to various incremental tasks, e.g., semantic segmentation [164, 165, 166, 167], person re-identification [168, 169], human action recognition [95], and federated learning [137]. However, since the knowledge distillation term aims to strike a balance between learning the new and remembering the old, it will undoubtedly suffer catastrophic forgetting compared to dynamic networks. [22] finds that knowledge distillation-based methods and dynamic networks have their advantages given different memory budgets. Specifically, knowledge distillation methods show stronger performance given limited memory, while dynamic networks require an adequate memory budget to perform competitively.
@@ 3.3.3 Discussions about Algorithm-Centric Methods
On the other hand, model rectification methods aim to reduce the biased inductive bias in the CIL model and align it to the oracle model. This line of work helps to understand the inherent factor of catastrophic forgetting. Apart from the rectification methods listed in Section 3.3.2, some methods [170, 116] address that the bias of the CIL model is from the imbalanced data stream. [171] finds that the embedding trained with contrastive loss suffers less forgetting than cross-entropy loss. [172] finds that the batch normalization layers [173] are biased in CIL and proposes to re-normalize the layer outputs. It is worth exploring other factors that could result in catastrophic forgetting and corresponding solutions in the future.
3.4 Discussions about Classification Criterion and Scopes of This Survey
In this paper, we divide current class-incremental learning methods into three groups based on the key point of these methods. For the methods discussing how to save and utilize exemplars, we divide them into data-centric methods. Similarly, for the methods designing model expansion and parameter-wise regularization, we denote them as model-centric methods. Lastly, we assign the methods that develop specific regularization to pass knowledge from the old to the new model and rectify the current model to the algorithm-centric group. It must be noted that these methods also learn from each other, and there is no strict boundary between them. For example, iCaRL [32] utilizes the knowledge distillation and exemplar replay, but we assign it to the knowledge distillation-based methods. In fact, data replay is becoming a standard protocol in the class-incremental learning community, which is widely adopted in most methods. As shown in Figure 3, dynamic networks and knowledge distillation methods have dominated publications in recent years. However, it does not indicate that traditional methods like data replay are no longer popular. By contrast, most of these methods rely on the exemplar set to get better performance.
On the other hand, in the discussion of this paper, we mainly focus on the development of class-incremental learning in the image classification field. The main reason is that most influential works about CIL address the image classification problem. Furthermore, it has shown the potential to apply the techniques in classification into other tasks, e.g., semantic segmentation [164, 165, 166, 167, 174], object detection [175, 176, 177, 178, 179, 180, 181], video understanding [95, 133], medical surgery [182, 183], language model [184, 185, 186, 187], language-vision model [188, 189], etc. Hence, the scope of our discussion is mainly focused on image classification tasks.

### v2

@@ 3.3.3 Prompt Expansion
Recently, Vision Transformer (ViT) [19] has attracted the attention of the computer vision community, and many works tend to design CIL learners using ViT as the backbone. DyTox [18] is the first work to explore ViT in CIL, which finds that model expansion in ViT is much easier than in convolutional networks. In DyTox, only task tokens are expanded for each new task, which requires much less memory than saving the whole backbone. Similarly, L2P [69] and DualPrompt [70] explore how to build CIL learners with pre-trained ViT. They borrow ideas from Visual Prompt Tuning (VPT) [179] to incrementally finetune the model with prompts. In L2P, the pre-trained ViT is frozen during the learning process, and the model only optimizes the prompts to fit new patterns. The prompt pool is defined as: $\mathbf{P}=\left\{P_{1},P_{2},\cdots,P_{M}\right\}$, where $M$ is the total number of prompts, $P_{i}\in\mathbb{R}^{L_{p}\times d}$ is a single prompt with token length $L_{p}$ and the same embedding size $d$ as the instance embedding $\phi({\bf x})$. The prompts are organized as key-value pairs — each instance selects the most similar prompts in the prompt pool via KNN search. It obtains instance-specific predictions by adapting the input embeddings as: ${{\bf x}}_{p}=\left[P_{s_{1}};\cdots;P_{s_{N}};\phi({\bf x})\right],\quad 1\leq N\leq M$, where $P_{s_{j}}$ are the selected instance-specific prompts. The adapted embeddings are then fed into the self-attention layers [32] to obtain instance-specific representations. CODA-Prompt [71] extends the prompt search with the attention mechanism. Apart from pre-trained ViT, S-Prompt [72] utilizes the pre-trained language-vision model CLIP [180] for CIL, which simultaneously learns language prompts and visual prompts to boost representative embeddings. Apart from expanding prompts, other lightweight modules can also be dynamically expanded in CIL [181, 182, 183]
Discussions: Learning dynamic networks, especially backbone expansion methods, has achieved state-of-the-art performance in recent years [17, 15, 16, 68]. However, it often requires expandable memory budgets, which is unsuitable for incremental learning on edge devices. To tackle this problem, further model compression [16], decoupling [17], and pruning can be adopted to alleviate the memory budget. Additionally, training DER requires an individual backbone for each task and aggregates all historical backbones as the feature extractor. It implicitly results in an unfair comparison to other methods with a single backbone [17]. In this paper, we systematically investigate the fair comparison protocol between these dynamic networks and others in Section 4.4, 4.5. Additionally, expanding backbones ignores the semantic information across tasks, e.g., when the old task contains ‘tigers’ and the new task contains ‘zebras,’ features like ‘stripes’ will be extracted by multiple backbones, resulting in feature redundancy. Hence, analyzing the semantic relationship across tasks [88] can help detect the feature redundancy, and contrastive learning [184] can be adopted for generalizable features.
On the other hand, most prompt expansion methods rely on the pre-trained models as initialization. Without such generalizable backbones, lightweight model updating with prompts often fails [185]. However, a pre-trained model is not always available for some specific downstream tasks, e.g., face recognition and speech recognition. Therefore, how to get rid of the dependence on pre-trained models is essential for these methods in real-world applications. Additionally, these prompt expansion-based methods tend to select instance-specific prompts based on a batch of instances. This requirement also needs to be satisfied during inference for accurate prompt retrieval [69], which implicitly results in an unfair comparison. Since a batch of instances is utilized to get the prompt, the context among instances becomes available, which is against the common sense of i.i.d. testing in machine learning.
Apart from these groups, there are works addressing network masks to divide a large network into sub-networks for each task [135, 132, 136]. However, deciding the activation of a specific sub-network requires the task identifier or learning extra task classifiers. On the other hand, several works [186, 152] propose to design specific modules for incremental new tasks, e.g., adapters [187]. However, manually handcrafting these modules requires heuristic designs or task-specific priors.
3.4 Parameter Regularization
@@
Dynamic networks seek to adjust model capacity with data evolves. However, if the model structure is fixed and unchangeable, how can we adjust the plasticity to resist catastrophic forgetting? Parameter regularization methods consider that the contribution of each parameter to the task is not equal. Hence, they seek to evaluate each parameter’s importance to the network and keep the important ones static to maintain former knowledge.
@@
The parameter-wise regularization term is calculated based on two parts. $\theta_{k}^{b-1}$ denotes the $k$-th parameter after learning last task $\mathcal{D}^{b-1}$. Hence, $(\theta_{k}^{b-1}-\theta_{k})^{2}$ represents the parameter drift from the last stage, and $\Omega_{k}$ weighs it to ensure important parameters do not shift away from the last stage. Since the model at the last stage represents the ‘old’ knowledge, consolidating important parameters can prevent the knowledge from being forgotten.
@@
Eq. 11 depicts a way to penalize essential parameters, and there are different ways to calculate the importance matrix $\Omega$. In EWC, Fisher information matrix [188] is adopted to estimate $\Omega$. However, the importance calculation in EWC is conducted at the end of each task, which ignores the optimization dynamics along the model training trajectory. To this end, SI [74] proposes to estimate $\Omega$ in an online manner and weigh the importance via its contribution to loss decay. RWalk [33] combines these importance estimation techniques. [75, 76] resort to an extra unlabeled dataset for online evaluation. IMM [77] finds a maximum of the mixture of Gaussian posteriors with the estimated Fisher information matrix. IADM [78] and CE-IDM [79] analyze the capacity and sustainability of different layers and find that different layers have different characteristics in CIL. In detail, shallow layers converge faster but have limited representation ability. By contrast, deep layers converge slowly while having powerful discrimination abilities. Hence, IADM augments EWC with an ensemble of different layers and learns layer-wise importance matrix in an online manner. K-FAC [80] extends the Fisher information matrix approximation with the Kronecker factorization technique.
Discussions: Although parameter and data regularization (Section 3.2) both exert regularization terms to resist forgetting, their basic idea differs substantially. Specifically, data regularization relies on the exemplar set to direct the optimization direction, while parameter regularization is based on the parameter-wise importance to construct the regularization term.
@@
As shown in Figure 3, parameter regularization methods have attracted the attention of the community in the early years. [189] shows that despite stemming from very different motivations, both SI [74] and MAS [75] approximate the square root of the Fisher Information, with the Fisher being the theoretically justified basis of EWC. However, estimating parameter importance requires saving the matrix with the same scale as the backbone. It faces the same risk as dynamic networks in that the memory budget is linearly increasing when learning more and more tasks.
On the other hand, the importance matrix may conflict at different incremental stages [88], making it hard to optimize the model and achieve poor performance on new tasks. Hence, although these works achieve competitive results in task-incremental learning, many works [8] find that parameter regularization-based methods perform poorly in the class-incremental learning scenario. To this end, some works try to alleviate the intransigence by learning a new backbone and consolidating them into a single one [77, 64]. In that case, the learning of new tasks will not be affected by the regularization term, and the parameter importance will only be considered during the consolidation process, enabling the model to be fully fitted to new tasks.
3.5 Knowledge Distillation
@@
Training data is evolving in the learning process, requiring tuning the model sequentially. We can denote the model after the previous stage $f^{b-1}$ as the ‘old model’ and the current updating model $f$ a s the ‘new model.’ Assuming the old model is a good classifier for all the seen classes in $\mathcal{Y}_{b-1}$, how can we utilize it to resist forgetting in the new model? To enable the old model to assist the new model, an intuitive way utilizes the concept proposed in [178], i.e., knowledge distillation (KD). KD enables the knowledge transfer from a teacher model to the student model, with which we can teach the new model not to forget. There are several ways to build the distillation relationship, and we divide these KD-based methods into three subgroups, i.e., logit distillation, feature distillation, and relational distillation.
@@
LwF [81] is the first success to apply knowledge distillation into CIL. Similar to Eq. 11, it builds the regularization term via knowledge distillation to resist forgetting:
@@
where the old model ${f}^{b-1}$ is frozen during updating. The regularization term builds the mapping between the old and new models by forcing the predicted probability among old classes to be the same. Given a specific input $\bf x$}, the output probability of the $k$-th class reveals the semantic similarity of the input to this class. Hence, Eq. 12 forces the semantic relationship of the old and new models to be the same and resists forgetting. iCaRL [82] extends LwF with the exemplar set, which helps to further recall former knowledge during incremental learning. Additionally, it drops the fully-connected layers and follows [190] to utilize nearest-mean-of-exemplars during inference. Eq. 12 strikes a trade-off between old and new classes, where the former part aims to learn new classes and the latter one maintains old knowledge. Since the number of old and new classes may differ in different incremental stages, BiC [83] extends Eq 12 by introducing a dynamic trade-off term:
@@
Figure 5: Illustration of knowledge distillation in CIL. Left: Logit distillation aligns the model outputs to make the old and new models share the same semantic relationship. Middle: Feature distillation aligns the features produced by the old and new models to ensure the new model does not forget old features. Right: Relational distillation resorts to structural inputs, e.g., triples, and aligns the input relationship of the old and new model.
@@
Eq. 13 forces the features extracted by the new embedding module to be the same as the old one, which is a stronger regularization than Eq. 12. Several works follow it to utilize feature distillation in CIL [90, 91, 97, 98], while others address distilling other products. LwM [92] suggests penalizing the changes in classifiers’ attention maps to resist forgetting. AFC [93] conducts the distillation considering the importance of different feature maps. PODNet [94] minimizes the difference of the pooled intermediate features in the height and width directions instead of performing element-wise distillations. CVIC [191] decouples the distillation term into spatial and temporal features for video classification. DDE [95] distills the causal effect from the old training to preserve the old knowledge. GeoDL [96] conducts distillation based on the projection of two sets of features from old and new models.
@@
Discussions: Knowledge distillation is a general idea to build the mapping between a set of methods, which has been widely adopted in class-incremental learning with many formats (e.g., logits, features, relationships). Due to their flexibility, knowledge distillation-based methods have also been widely applied to various incremental learning tasks, e.g., semantic segmentation [193], person re-identification [194], human action recognition [91], and federated learning [149]. Since a set of models exists in CIL, it is intuitive to build the student-teacher mapping in CIL, making knowledge distillation an essential solution for most works.
However, since the knowledge distillation term aims to strike a balance between learning the new and remembering the old, it is hard to control the precise trade-off term between plasticity and stability. Specifically, giving more importance to the knowledge distillation term will harm the plasticity of learning new tasks, while giving lower importance will cause catastrophic forgetting or feature overwriting. Compared to dynamic networks, knowledge distillation-based methods lack the ability to learn more informative features as data evolves. [17] compares knowledge distillation-based and dynamic network-based methods and finds that knowledge distillation-based methods and dynamic networks have their advantages given different memory budgets. Specifically, knowledge distillation methods show stronger performance given limited memory, while dynamic networks require an adequate memory budget to perform competitively. Besides, feature/relational distillation only regularizes the extracted features to be similar, thus regularizing the embedding function to resist forgetting. However, due to the data characteristics of CIL, the classifier layer also gets biased and forgets former knowledge and knowledge distillation-based methods cannot handle such challenges.
3.6 Model Rectify
@@
The first method addressing the bias of the CIL model is UCIR [89]. It finds that the weight norm of new classes is significantly larger than old ones, and the model tends to predict instances as the new classes with larger weights. Hence, UCIR proposes to utilize a cosine classifier to avoid the influence of biased classifiers: $f({\bf x})=\frac{W}{\lVert W\rVert}^{\top}\frac{\phi({\bf x})}{\lVert\phi({\bf x})\rVert}$. Thus, the weight norm will not influence model predictions in incremental learning. WA [112] further normalizes the weight after every optimization step. It also introduces weight clipping to ensure the predicted probability is proportionate to classifier weights. SS-IL [113] explains the reason for weight drifting, which is caused by the imbalance phenomena between old and new instances. Since the number of new class instances is much more than that of old ones, optimizing the model with cross-entropy loss will increase the weight of new classes and decrease old ones. Hence, SS-IL suggests separated softmax operation and task-wise knowledge distillation to alleviate the influence of imbalanced data. RPC [114] claims that the classifiers of all classes can be pre-allocated and fixed. This makes it impossible for classifiers to be biased toward new classes.
@@
Only the logits for new classes ($k\in{Y}_{b}$) are rectified after each incremental task. BiC separates an extra validation set from the exemplar set, i.e., $\mathcal{E}=\mathcal{E}_{train}\cup\mathcal{E}_{val}$, and uses the validation set to tune the rectification layer. On the other hand, IL2M [111] suggests re-scaling the outputs with historical statistics. Suppose an instance is predicted as a new class. In that case, the logits will be re-scaled to ensure the predictions of old and new classes follow the same distribution.
@@
Lastly, since the embedding module is sequentially updated in CIL, some works try to rectify the biased representations of incremental models. For example, SDC [105] utilizes a nearest-mean-of-exemplars classifier, which calculates the class centers and assigns instances to the nearest class center. However, since the embedding is incrementally updated, the class centers calculated in the former stage may suffer a drift in the next stage, making the classification results unreliable. Since old class instances are not available in the current stage, SDC aims to calibrate the class centers of old classes with the drift of new classes. CwD [106] analyzes the differences of embeddings among the CIL model and the oracle model and finds that the embeddings of the oracle model scatter more uniformly. It aims to make the CIL model similar to the oracle by enforcing the eigenvalues to be close. ConFiT [107] relieves the feature drift from the middle layers. MRFA [195] finds the all-layer margin of replay samples shrinks as data evolves. Other works address the model weight rectification. CCLL [108] aims to calibrate the activation maps of old models during incremental learning. RKR [109] proposes to rectify the convolutional weights of old models when learning new tasks. FACT [104] depicts a new training paradigm for CIL, namely forward compatible training. Since the embedding space is endlessly adjusted for new classes, FACT proposes to pre-assign the embedding space for new classes to relieve the burden of embedding tuning.
@@
Discussions: Model rectify-based methods aim to reduce the inductive bias in the CIL model and align it to the oracle model. This line of work helps to understand the inherent factor of catastrophic forgetting. Apart from the rectification methods listed in this section, [113] addresses that the bias of the CIL model is from the imbalanced data stream. [184] finds that the embedding trained with contrastive loss suffers less forgetting than cross-entropy loss. [196] finds that the batch normalization layers [197] are biased in CIL and proposes re-normalizing the layer outputs. [198] finds that vision transformers gradually lose the locality information when incrementally updated and proposes to insert the prior information about locality into the self-attention process. These works often treat the oracle model as the example, and design training techniques to reflect oracle model’s characteristics.
Apart from mimicking the oracle model, there are also works addressing the forward compatibility [104] and flat loss landscape. Since the ultimate goal of CIL is to find a flat minimum in loss landscape among all tasks, several works aim to achieve this goal during the first stage [199, 200]. It is worth exploring other factors in catastrophic forgetting and corresponding solutions in the future. On the other hand, the oracle model is obtained via joint training of all data via supervised loss. Other task-agnostic features could also help build a holistic classifier, e.g., via contrastive learning [184], which the oracle model does not possess.
3.7 Template-Based Classification
Lastly, we discuss template-based classification, which is widely adopted in CIL. If we can build a ‘template’ for each class, the classification can be done by matching the query instance to the most similar template. A popular approach is to utilize class prototypes [129] as the template, which is rooted in cognitive science [201]. The prototype in deep neural networks is often defined as the average vector in the embedding space. For example, we can utilize the current embedding function $\phi(\cdot)$ to extract the prototype of the $i$-th class:
${\bm{p}}_{i}=\frac{1}{N}\sum_{j=1}^{|\mathcal{{D}}^{b}|}\mathbb{I}(y_{j}=i)\phi({\bf x}_{j})\,$, (16)
where $N$ is the instance number of class $i$. In Eq. 16, the class prototype is calculated via the class center in the embedding space. Hence, we can make inference without relying on the fully connected layer by matching an instance to the nearest prototype:
$y^{*}=\underset{y=1,\ldots,|\mathcal{Y}_{b}|}{\operatorname{argmin}}\left\|\phi({\bf x})-{\bm{p}}_{y}\right\|\,.$ (17)
As discussed in model rectification-based methods, sequentially updating the incremental model will result in bias in the fully-connected layer. Correspondingly, iCaRL [82] suggests conducting inferences via Eq. 17, which is also known as nearest-class-mean classifier [190]. Since inference is conducted by instance-prototype matching in the same embedding space, the bias among different stages can be alleviated. However, utilizing prototype-based inference also faces another challenge, i.e., the embedding mismatch between stages. Since the embedding function keeps changing among different stages, the prototypes of former stages may be incompatible with the query embedding of later stages. This phenomenon is also known as the semantic drift [105]. To fill in this gap, a naive solution is to utilize the set of exemplars and re-calculate class prototypes after each stage [82, 37]. Since the exemplar set $\mathcal{E}$ contains instances of former classes, re-calculating all class prototypes after each incremental stage ensures the compatibility between prototypes and the latest embedding function. However, when exemplars are unavailable, specific algorithms need to be designed to compensate for the semantic drift.
Prototype-based Inference without Exemplars: When exemplars are unavailable, there are two main solutions to maintain a prototype classifier. A naive solution is to freeze the embedding function after the first incremental stage [104, 116], which forces the prototypes of different stages to be compatible. Such a learning process assumes the embedding function trained with first-stage data is generalizable enough for future tasks, relying on the vast number of training instances in $\mathcal{D}^{1}$. As a result, researchers tend to design suitable training techniques with $\mathcal{D}^{1}$ to obtain a generalizable feature space for future tasks. [120, 49] draw inspiration from contrastive learning and design pre-text tasks to enhance the representation learned by the first stage. CEC [115] meta-learns a graph model to adjust new class prototypes with known ones, which propagates context information between classifiers for adaptation. LIMIT [121] finds that using prototypes extracted by the first stage backbone tends to predict instances into classes of the first stage. It proposes calibrating the prediction logits by meta-learning a transformer block between old and new classes. TEEN [119] systematically analyzes the performance gap between old and new classes and finds that the prototypical network forces the model to predict new classes into the most similar old class. Hence, it suggests pushing the prototypes of new classes to old classes for a calibrated decision boundary. Moreover, [104] aims to enhance the model’s forward compatibility by reserving the embedding space for new classes so that new classes can be inserted into the embedding space without harming existing ones. It allocates ‘virtual prototypes’ for new classes and explicitly reserves the embedding space for them using a bimodal target label and manifold mixup [202] to generate new class instances.
When we have a pre-trained embedding function as initialization, ADAM [116] finds that using a prototype-based classifier easily beats the state-of-the-art prompt-based methods [70, 69]. However, although the pre-trained model possesses generalizable features, it still lacks task-specific information on incremental datasets. Hence, it designs the ‘adapt and merge’ protocol to unify the generalizability of the pre-trained model and the adaptivity of downstream tasks. With the pre-trained embedding, it finetunes it with the first stage dataset $\mathcal{D}^{1}$, and concatenates the pre-trained and adapted embeddings for prototype extraction. Based on ADAM, RanPAC [117] designs random projection to project the concatenated features into a high-dimensional space, within which classes are separated more clearly. It also incrementally updates a Mahalanobis distance-based classifier, which shows stronger performance than cosine distance between prototypes and query embedding. FeCAM [118] also finds the inadequacy of the cosine classifier and proposes using a Bayesian classifier instead.
Apart from freezing the embedding, some works try to compensate for the semantic drift among different stages. Since prototypes will drift with the ever-changing embedding functions, they aim to estimate such drift to estimate the prototypes in the latest embedding space. SDC [105] aims to measure the prototype drift between different stages and utilizes the weighted combination of current stage data for reference. ZSTCI [122] achieves this goal by mapping prototypes of different stages into the same embedding space and designing a prototype alignment loss across stages.
Finally, another line of work considers generative classification [123]. Different from estimating the class prototype as a template, the template for each class is a generative model. Hence, the inference process can be measured via the likelihood of the query instance under such a generative model. However, it requires more calculation budgets for generative templates during inference than prototype-based methods.
Discussions: There are two advantages of template-based classification. Firstly, when the exemplar set is available, using template-based classification enables query-prototype matching in the embedding space. Since the classifier will be biased after incremental learning, utilizing such a matching target alleviates the inductive bias during inference [82]. Secondly, when the pre-trained model (or model trained with large base classes) is available as initialization, the feature representations are generalizable and can be transferred to downstream tasks. Hence, freezing the embedding and using template-based classification can take full use of the generalizable features, and such a non-incremental learner will not suffer forgetting due to the frozen backbone [116, 117, 118].
However, there are also some drawbacks. When the exemplar set is not available, re-calculating the prototypes is impossible, and it requires a complex adjusting process to overcome the semantic drift [105, 122]. Secondly, when freezing the backbone and taking a template-based classifier, the model sacrifices its adaptivity for downstream tasks. When there are significant domain gaps between the pre-trained model and the downstream data [203, 204], template-based classification shall fail due to the incapability to extract generalizable features. In that case, continually adjusting the backbone could be more suitable to extract task-specific features. Finally, we can use an energy-based model [205] to compute for each class an energy value rather than a likelihood [206] to alleviate the high cost of the generative model.